# Reproducible Machine Learning-based Voice Pathology Detection: Introducing the Pitch Difference Feature

### v2

@@ 9 Conclusion
  • we provided reliable and reproducible results showing the top performance for selected classifiers, reporting realistic performance values,
  • we introduced two novel features, pitch difference and NaN feature, which both were represented in the feature subsets that reached the best performance for a given classifier,
@@
  • we presented the top performing feature sets for each classifier type with respect to all tested hyperparameter configurations.
Finally, there are several limitations our work which are mostly based on lack of suitable databases. Due to this fact, we omitted the external validation of reached results.
Funding sources
@@
CRediT Authorship Contribution Statement
@@
Declaration of Competing Interest
Supplementary Materials
list_of_excluded_filescsv:
- https://doi.org/10.5281/zenodo.14793017
@@ Appendix A Reforms checklist
In our research, we examine the feasibility of ML methods for voice pathology detection in adult patients. The SVD is the only available suitable dataset, having a relatively large number of samples, while containing a wide number of diseases, which are also presented in the general population (see Section 2 for reasoning and Section 3 for description of the data).
@@
We utilize the publicly available dataset "Saarbruecken Voice Database" available at: https://stimmdb.coli.uni-saarland.de/help_en.php4. There is no unique identifier and no version control of the database, and we do not have a license to share the dataset, therefore, we provide list of files along with sha256 checksum to ensure reproducibility (see Section 3).
@@
All data are from Saarbruecken Voice Database as described in 2a. As this database is relatively small, we utilize only stratified 10-fold cross-validation without test or evaluation dataset. We downloaded data from SVD on July 12, 2022. The recordings used in our study were recorded between November 20, 1997 and June 16, 2004. The ground truth annotations were obtained by evaluation of stroboscopical recording by database authors [28]. All information related to SVD can be found at https://stimmdb.coli.uni-saarland.de/help_en.php4.
@@
We believe SVD dataset is relevant as it contains various pathologies that are also present in general population. In our study, after removing underaged patients and duplicities, we worked with 64 various pathologies (see Section 3).
@@
The outcome variable is derived from the table containing patient information. Classification takes the value of 0 if there are no values in the pathology column in the svd_information.csv file, and the value of 1 if one or more pathologies are listed in the column (see Section 3). The distribution of healthy / pathological recordings of female and male subjects is provided in the Table 1.
@@
In our study, we utilized 1636 recordings out of 2041 total. In total, 977 are pathological and 659 are healthy. More detailed distribution is explained in Table 1, Figure 4, and Figure 5 (see Section 3).
@@
There was no missing data. However, 1 file (1573-a-n.wav) contains a recording of the time between two sessions and one is corrupted (87-a-n.wav). Moreover, there were multiple recordings marked as corrupt in comments to recording sessions. These recordings were excluded from our research. Finally, there were multiple recordings marked with pathologies "Gesangsstimme" and "Sängerstimme", which probably contain healthy subjects. However, as it is not clear, we exclude this data unless they are diagnosed with another pathology (see Section 3).
@@
From the dataset, we remove 13 corrupted recordings, 40 underage recordings, 17 singers and 335 recordings of patients with multiple recording sessions (except first healthy and first pathological recordings, see Figure 5). See Section 3 for rationale and GitHub repository for list of excluded files (misc/ list_of_excluded_files.csv).
@@
When the extraction of f0 was impossible, the feature values for $\overline{f}_{0}$, $\sigma_{\mathbf{f}_{0}}$, $\Delta f_{0}$, jitta, and shim were set to 0 and the binary NaN feature was set to 1 (see Section 4).
@@
See Figure 5. More details are in sections 3, 4, and 5.
@@
We utilize multiple ML algorithms for classification (see Table 3) and the k-means SMOTE algorithm for dataset augmentation (see Section 5).
@@
All ML algorithms are suitable for multi-dimensional data, that we are dealing with (see Section 6).
@@
Information about stratified 10-fold cross-validation and repeated stratified 10-fold cross-validation for the best models is described in the section Section 7.
@@
We performed repeated 10-fold cross-validation to estimate the average value of MCC and its corresponding standard deviation. We select the best model according to this MCC. See Section 7.
@@
By applying the oversampling algorithm only to the training folds, we aimed to prevent data leakage and ensure that the model’s performance evaluation on the test fold remains unbiased. This approach allowed us to mitigate the adverse effects of class imbalance, resulting in a more robust and reliable predictive model. See Sections 5 and 7.
@@
For patients with multiple recordings of the same type (either all healthy or all pathological), we retained only the oldest recorded sample. For patients with both healthy and pathological recordings, we selected the oldest sample of each type, resulting in a maximum of two recordings per patient — one healthy and one pathological. We believe this approach minimizes the likelihood of the classifier learning patient identities, as the patient’s classification remains independent of their identity. See Section 3.
@@
All features, except two, are obtained from voice recordings and are widely used in the models for voice pathology detection (see Section 2). We consider the "AGE" feature legitimate, as other acoustic features depend on the AGE. I.e. there are changes in speaking fundamental frequency with aging [53].
The feature, that we introduced as "NaN" reflects the fact, that it was not possible to estimate the fundamental frequency for the patient. As the fundamental frequency is considered as one of the dominant features in voice pathology diagnosis, we consider introducing this "NaN" feature a legitimate approach. Note, that in total, there are 2 females and 6 males in our dataset, that have NaN value of fundamental frequency, all of them are suffering the voice disorder (see Section 4).
@@
The choice of metrics, with respect to the class imbalance in the data, is written in the Section 7. The claim regarding the best model is based on the Matthews correlation coefficient metric, that is suitable for imbalanced datasets and reflect both successes and errors in the classification.
@@
For each of metrics specified in Section 7, we provide also the respective standard deviations that were obtained during the cross-validation procedure which is specified in this section.
@@
First, our model was tested using SVD only. The used database does not fully reflect the general population, in the sense of proportion of healthy / pathological voices. However, at this time, there is no other suitable database that would reflect the general population better. Despite the justification as an only viable source of data, we cannot extrapolate its performance outside of this dataset. Moreover, we limit our research only to individuals that are 18 years old and older. As the data was recorded in a controlled environment, we can assume our classifiers might not be able to perform as well with datasets that are recorded during different conditions. Another noteworthy limitation was the available computational capacity which led to careful decision of the classification algorithms and hyperparameter space we drew from during our work. See Section 8.
Appendix B Features used
B.1 Time-Domain Features
The majority of the time-domain features stem from the measurement of regularity of the sound signal. The idea behind the measurement of (ir)regularities in a voice recording is that human voice is composed of periodic and random components and there is a difference between the amount of irregularity between the healthy and pathological voices.
B.1.1 Fundamental frequency
The periodic part is characterized by f0. Voice pathologies usually increase the ratio between the random and periodic components, also altering f0. Thus, pathologies might be observable from the change of these parameters.
The main disadvantage of several time-domain measures is the necessity to estimate f0. In heavily altered voices, caused by some pathologies, it might be impossible to estimate f0. There are also multiple algorithms for f0 estimation and generally, each may yield different results.
We used the parselmouth library, a Python implementation of the Praat program [54], to determine f0, therefore, we utilized the algorithm described in [55] to extract f0 for $N$ segments of the samples, denoted as $\mathbf{f}_{0}$. Given that the estimation of f0 is calculated for individual segments of each recording, we took the mean (Equation 8) and sample standard deviation (Equation 9) of f0 as features for each sample.
$\overline{f}_{0}=\cfrac{1}{N}\sum_{i=0}^{N-1}f_{0,i}$ (8)
$\sigma_{\mathbf{f}_{0}}=\sqrt{\cfrac{\sum_{i=0}^{N-1}(f_{0,i}-\overline{f}_{0})^{2}}{N-1}}$ (9)
B.1.2 Jitter
Jitter, also called fundamental frequency perturbation, is a measure of irregularities in f0. Jitter is a well-known metric, often used in acoustic analysis for different types of signals. In voice pathology, higher values of jitter usually indicate pathology.
There are several variations of the measure, in our work, we used the local jitter, also known as jitta (Equation 10) [54], calculated as the mean difference between periods of consecutive glottal cycles $T_{i}$ across $N$ glottal cycles divided by the mean glottal cycle period of the signal $\overline{T}$ (Equation 11).
$\text{jitta}=\cfrac{\sum_{i=1}^{N-1}\lvert T_{i}-T_{i-1}\rvert}{(N-1)\cdot\overline{T}}$ (10)
$\overline{T}=\sum_{i=0}^{N-1}\cfrac{T_{i}}{N}$ (11)
B.1.3 Shimmer
Shimmer, also known as amplitude perturbation, measures irregularities in the amplitude of the sound wave and high values of shimmer also indicate potential voice pathology.
There are several variations of shimmer similar to the jitter and we opted for the local shimmer, also called shim (Equation 12) [54]. Shim is calculated as the mean difference between amplitudes of consecutive glottal cycles $A_{i}$ across $N$ glottal cycles divided by the mean glottal cycle amplitude of the signal $\overline{A}$ (Equation 13).
$\text{shim}=\cfrac{\sum_{i=1}^{N-1}\lvert A_{i}-A_{i-1}\rvert}{(N-1)\cdot\overline{A}}$ (12)
$\overline{A}=\sum_{i=0}^{N-1}\cfrac{A_{i}}{N}$ (13)
B.1.4 Harmonics-to-Noise Ratio
HNR measures the ratio between the energies of the periodic and aperiodic signal components (Equation 14) [55, 54].
$\text{HNR}=10\log_{10}\cfrac{\mathcal{X}_{i_{max}}}{1-\mathcal{X}_{i_{max}}}$ (14)
We use the recommended cross-correlation implementation of the harmonicity algorithm where the energy of the harmonic component in a windowed signal $\mathcal{X}_{k_{max}}$ is calculated as the value of a cross-correlation of the signal amplitude $s$ with the same signal shifted by the length $k_{max}$ where the cross-correlation yields the highest result (Equation 15). Note that the energy is normalized by the cross-correlation of the signal with itself without any shift.
$\mathcal{X}_{i,k}=\sum_{j=m}^{m+n-1}s_{j}s_{j+k}$ (15)
In Equation 15, $i$ represents the window index, $k$ represents the index of a value in the $i$-th window, $m$ represents the index of an element with respect to the current window, and $n$ represents the number of samples in a window. The length of the window is chosen with respect to the expected f0.
B.1.5 Zero-Crossing Rate
Zero-crossing rate (ZCR) measures the rate at which a sound signal $s$ changes signs in a signal split to a predetermined number of windows. This feature is not dependent on the estimation of f0. The calculation for the $i$-th window is described in Equation 16, where $k$ represents the index of a value in the $i$-th window, $m$ represents the index of an element with respect to the current window, and $n$ represents the number of samples in a window, set to default value of 2048. Note that $\mathbb{1}$ is described in Equation 17, and $\text{sgn}(x)$ is described in Equation 18.
$\text{ZCR}_{i}=\cfrac{1}{n-1}\sum_{j=m}^{m+n-1}\mathbb{1}(s_{j},s_{j+1})$ (16)
$\mathbb{1}(s_{j},s_{j+1})=\begin{cases}1&\text{sgn}{(s_{j})}\,\text{sgn}{(s_{j+1})}=-1\\ 0&\text{sgn}{(s_{j})}\,\text{sgn}{(s_{j+1})}=+1\end{cases}\,\,\,$ (17)
$\text{sgn}{(x)}=\begin{cases}-1&x<0\\ +1&x\geq 0\end{cases}$ (18)
B.1.6 Shannon Entropy
Shannon entropy measures the average level of uncertainty inferent to the possible outcomes of a random variable. Generally, it is defined as in Equation 19 for an independent variable $X$ defined by a set of values $\mathcal{X}$ with a probability of occurrence $p$.
$H(X)=-\sum_{x\in{\mathcal{X}}}p(x)\log p(x)$ (19)
In our case, the probabilities $p$ are taken from the histogram of the signal amplitude. The number of bins $n_{bins}$ is determined by the Freedman–Diaconis rule (Equation 20), where $N$ is the length of the signal and $Q_{3}$, $Q_{1}$ are amplitude values representing the third and first quartiles, respectively.
$n_{bins}=2\cdot\cfrac{Q_{3}-Q_{1}}{\sqrt[3]{N}}$ (20)
This feature is also independent of f0.
B.1.7 Skewness
Skewness measures the concentration of values on either side of a data distribution. Based on the results of the study [40], we expected that the skewness value might be influenced by a pathology, and thus, we included this feature in our experiments.
We calculate the skewness according to the definition in Equation 21.
$\text{skew}=\cfrac{m_{3}}{m_{2}^{3/2}}$ (21)
The individual cumulants $m_{i}$ can be determined by Equation 22.
$m_{i}=\cfrac{1}{N}\sum_{j=1}^{N}(s_{j}-\overline{s})^{i}$ (22)
B.2 Spectral Features
We use spectral features to measure characteristics and differences in the spectra of the sound signal, as this is a common approach in voice pathology detection (i.e. [41, 56, 42]). The most frequent transformation for obtaining spectral information is the Fourier transform. To gain more information about the spectral variations of the signal in time, the signal is split into short windows containing tens to hundreds of milliseconds of the original signal with varying overlaps between these windows. Then, the short-time Fourier transform (STFT) algorithm is applied to each window to gain the spectrum of each window. Finally, the spectra are concatenated and plotted in the form of a spectrogram - a heatmap showing time on the x-axis, frequency on the y-axis, and power of the spectrum as a color. In regard to the voice pathology detection, the expectations are that due to the difference in the ratio of (dis)harmonic components, the spectral features should be able to distinguish between pathological and healthy voices.
B.2.1 Spectral Flatness
Spectral flatness is used to quantify how much tone-like a signal is, as opposed to being noise-like [57]. It is calculated as a ratio of the geometric mean of the power spectrum $S$ to its arithmetic mean measured across all bands [58] (see Equation 23), where $N$ denotes the total number of bands of the power spectrum.
$\text{flatness}=\cfrac{\sqrt[N]{\prod_{n=1}^{N}S(n)}}{\cfrac{1}{N}\sum_{n=1}^{N}S(n)}=\cfrac{\exp\Bigg{(}\cfrac{1}{N}\sum_{n=1}^{N}\ln{S(n)}\Bigg{)}}{\cfrac{1}{N}\sum_{n=1}^{N}S(n)}$ (23)
In this form, spectral flatness reaches values from 00, representing a pure tone, to $1$, representing a white noise with equal energy distribution across all bands. We use the representation from [57, 58], implemented in the librosa library [33] for the purpose of our work. Due to practical implementation, the zero values of the power spectrum are replaced by the value of $10^{-10}$.
B.2.2 Spectral Roll-Off
Spectral roll-off shows the frequency band such that at least a predetermined part of the spectral energy is contained in this and previous bins. We used the threshold value of $85\%$ during our experiments.
B.2.3 Mean Spectral Centroid
Spectral centroid is the simple spectral shape feature, that is used in some works on voice pathology detection. Assuming the normalized spectrum $\tilde{S}_{i}$ in the $i$-th time frame and the set of frequency bins $K$ given as
$\tilde{S}_{i}(f)=\cfrac{\lvert S_{i}(f)\rvert}{\sum_{j\in K}\lvert S_{i}(j)\rvert}\ \ ,$ (24)
we can compute the spectral centroid $C_{i}$ as
$C_{i}=\sum_{f\in K}f\tilde{S_{i}}(f)\ .$ (25)
To get the single-value feature, and avoid the problem with different lengths of recording, we used the mean value of spectral centroid as a feature.
B.2.4 Spectral Contrast
Spectral contrast measures the difference between the mean energy of the top and bottom quantiles of selected energy bands. We calculate the spectral contrast with the librosa library, which implements it as described in Algorithm 2.
Algorithm 2 Spectral contrast
1:create a spectrogram using STFT with Hann window of length 2024
2:divide the frequency domain into 7 bands octave-scale sub-bands, 0 - 200 Hz, 200 - 400 Hz, 400 - 800 Hz, 800 - 1600 Hz, 1600 - 3200 Hz, 3200 - 6400 Hz, 6400 - 11025 Hz (recordings are resampled to 22050 Hz)
3:for each band $k$ and time window $i$ of length 2048 do
4:     order the values for each frequency band in each time window in ascending order
5:     get the $\text{peak}_{k,i}$ value as a mean value from first $n$ values, where $n$ represents the number of values in the top 2%
6:     get the $\text{valley}_{k,i}$ value as a mean value from last $n$ values, where $n$ represents the number of values in the bottom 2%
7:     compute spectral contrast of $k$th band and $i$th time windows as $\text{SC}_{i,k}=\text{peak}_{i,k}-\text{valley}_{i,k}$
8:end for
9:Calculate the average spectral contrast for each band $\text{SC}_{mean,k}$ (Equation 26), where $N$ represents the total number of time windows.
$\text{SC}_{mean,k}=\cfrac{1}{N}\sum_{i=0}^{N-1}\mathbf{SC}_{i,k}$ (26)
B.2.5 Formants
Formants characterize peaks in frequency spectra of a human voice. In linguistics, they are often used to differentiate between different vowels. The most prominent are the first two formants, based on which it is usually possible to distinguish between the vowels [59].
We extracted the frequencies of the first three formants using the parselmouth library.
B.3 Cepstral Features
While time-domain and spectral features are common in many signal processing fields, the cepstral features are more prominent in the voice analysis. Cepstral features are derived from a cepstrum, which describes the spectral information the same way spectrum describes the time-domain information of the signal.
B.3.1 Mel-Frequency Cepstral Coefficients
Cepstral coefficients are used in voice and musical analysis to highlight differences in the power between individual frequencies of a sound signal.
The construction of a cepstrogram involves several steps. We explain the algorithms implemented in the librosa library we used for the feature extraction.
First, we take a power spectrogram of the recording, using the STFT algorithm defined in Equation 27. $X(k,f)$ represents the transformation as a function of time frame $k$ and frequency band $f$. The algorithm utilizes a window $w_{n-k}$, in our case, we use a Hann window centered around time frame $k$ with the size $N=2048$ samples and overlap of $N/4$ or $75\%$, which are default values of the algorithm.
$X(k,f)=\sum_{n=-\infty}^{+\infty}x_{n}\cdot w_{n-k}\cdot\exp{(-2\pi jfn)}$ (27)
Next, we transform $X$ into a power spectrogram $P[dB]$ as in Equation 28. In the equation, we describe the transformation of each value in the time frame $k$ and frequency band $f$.
$P[\mathrm{dB}]_{k,f}=10\log_{10}(\lvert X(k,f)\rvert^{2})$ (28)
Next, the power spectrogram $P[\mathrm{dB}]$ is transformed to a mel spectrogram $P[\mathrm{mel}]$ using a mel filterbank $H$. The mel filterbank represents a series of triangular filters with their centers evenly distributed on the mel scale. We choose the default number of filters for the filterbank $M=128$. The centers would be evenly distributed between $f_{\text{mel, min}}=0$ and $f_{\text{mel, max}}=3176$, which represents the maximum frequency of $11025$ Hz, half of the sampling frequency. The maximum frequency is chosen with respect to the Nyquist theorem. The formula for transforming hertz frequency to mel scale is in Equation 29 and it can be transformed back by Equation 30.
$f_{\text{mel}}=2595\log_{10}\biggl{(}1+\cfrac{f}{700}\biggr{)}$ (29)
$f=700\biggl{(}10^{\bigl{(}\frac{f_{\text{mel}}}{2595}\bigr{)}}-1\biggr{)}$ (30)
The filterbank $H_{m,f}$ for the $m$-th filter and a frequency band $f$ is then calculated by Equation 31. Note that $f_{m}$ here represents the frequency center of the $m$-th filter in Hz and $f$ represents a frequency band in the time frame.
$H_{m,f}=\begin{cases}0&f<f_{m-1}\\ \cfrac{f-f_{m-1}}{f_{m}-f_{m-1}}&f_{m-1}\leq f<f_{m}\\ \cfrac{f_{m+1}-f}{f_{m+1}-f_{m}}&f_{m}\leq f<f_{m+1}\\ 0&f>f_{m+1}\\ \end{cases}$ (31)
Now, we can calculate the $m$-th band of the mel spectrogram $P[\mathrm{mel}]_{m}$ by multiplying each time frame $k$ of the spectrogram $P[dB]_{k}$ by the amplitude for each frequency band $f$ of the respective filter bank $H_{m,f}$ as in Equation 32. The resulting mel spectrogram has dimensions $[M,N]$, where $M$ is the number of filters in the mel filterbank and $N$ is the number of frames of the STFT.
$P[\mathrm{mel}]_{m}=\sum_{k=0}^{N-1}X_{k}H_{m,f}$ (32)
Finally, the cepstrogram or the MFCC are calculated by transforming the mel spectrogram with discrete cosine transform (DCT). The $l$-th band of the cepstrogram $C$ is calculated by Equation 33.
$C_{l}=2\sum_{m=0}^{M}P_{\text{mel}}[m]\cos{\biggl{(}\cfrac{\pi l(2m+1)}{2M}\biggr{)}}$ (33)
The result is a cepstrogram with dimensions $[L,N]$, where $L$ represents the number of taken MFCC. Since the coefficients are different across the time windows, we took the mean value of each band as well as the variance of the values in each band as our features.
Moreover, we also estimated the first and second derivatives from the extracted MFCC using the Savitzky-Golay filter. Similar to the coefficients, we took the mean values and variance of the derivatives as our features.
B.3.2 Linear-Frequency Cepstral Coefficients
LFCC are constructed similarly to the MFCC with one distinction, the triangular filters applied to the frequency spectrum are spaced equidistantly in the standard frequency scale.

### v3

@@ 6 Conclusion
  • 1.
    we provided reliable and reproducible results showing the top performance for models based on the selected ML algorithms and hyperparameter settings, reporting realistic performance values,
  • 2.
    we introduced two novel features, pitch difference and NaN feature, which both were represented in the feature subsets that reached the reported best performance,
  • 3.
@@
  • 5.
    we presented the top performing feature sets for each ML algorithm with respect to all tested hyperparameter configurations.
Finally, there are several limitations to our work which are mostly based on a lack of suitable databases. Due to this fact, we omitted the external validation of the reached results.
Supplementary information
list_of_excluded_files.csv: recordings excluded along with reasons for exclusion (as described in Section 3.1).
Acknowledgements
@@
Declarations
Data availability
All data used are publicly accessible on the following webpage: https://stimmdb.coli.uni-saarland.de
Code availability
The code used to produce all results, along with all supplemental material and information to reproduce our results, is stored in publicly available GitHub repository https://github.com/aailab-uct/Automated-Robust-and-Reproducible-Voice-Pathology-Detection with the following DOI: 10.5281/zenodo.13771573.
Author contribution
@@
Glossary
BM
bookmakers informedness
CNN
convolutional neural network
DNN
deep neural networks
DT
decision trees
EGG
electroglottogram
f0
fundamental frequency
GFCC
gammatone frequency cepstral coefficients
GM
geometric mean
HNR
harmonic-to-noise ratio
KNN
k-nearest neighbors
LFCC
linear-frequency cepstral coefficients
LPC
linear predictive coefficients
LSTM
long short-term memory
MCC
Matthews correlation coefficient
MFCC
mel-frequency cepstral coefficients
ML
machine learning
MLP
multi-layered perceptron
NB
naive Bayes
NN
neural network
PRNG
pseudo-random number generator
RF
random forest
SMOTE
synthetic minority over-sampling technique
SVD
Saarbrücken Voice Database
SVM
support vector machine
UAR
unweighted average recall
@@ Appendix A Reforms checklist
In our research, we examine the feasibility of ML methods for voice pathology detection in adult patients. The SVD is the only available suitable dataset, having a relatively large number of samples, while containing a wide number of diseases, which are also presented in the general population (see Section 2 for reasoning and Section 3.1 for description of the data).
@@
We utilize the publicly available dataset ”Saarbruecken Voice Database” available at: https://stimmdb.coli.uni-saarland.de/help_en.php4. There is no unique identifier and no version control of the database, and we do not have a license to share the dataset, therefore, we provide list of files along with sha256 checksum to ensure reproducibility (see Section 3.1).
@@
All data are from Saarbruecken Voice Database as described in 2a. As this database is relatively small, we utilize only stratified 10-fold cross-validation without test or evaluation dataset. We downloaded data from SVD on July 12, 2022. The recordings used in our study were recorded between November 20, 1997 and June 16, 2004. The ground truth annotations were obtained by evaluation of stroboscopical recording by database authors [36]. All information related to SVD can be found at https://stimmdb.coli.uni-saarland.de.
@@
We believe SVD dataset is relevant as it contains various pathologies that are also present in general population. In our study, after removing inappropriate recordings, we worked with 64 various pathologies (see Section 3.1).
@@
The outcome variable is derived from the table containing patient information. Classification takes the value of 0 if there are no values in the pathology column in the svd_information.csv file, and the value of 1 if one or more pathologies are listed in the column (see Section 3.1). The distribution of healthy / pathological recordings of female and male subjects is provided in the Table 1.
@@
In our study, we utilized 1636 recordings out of 2041 total. In total, 977 are pathological and 659 are healthy. More detailed distribution is explained in Table 1, Figure 4, and Figure 5 (see Section 3.1).
@@
There was no missing data. However, 1 file (1573-a-n.wav) contains a recording of the time between two sessions and one is corrupted (87-a-n.wav). Moreover, there were multiple recordings marked as corrupt in comments to recording sessions. These recordings were excluded from our research. Finally, there were multiple recordings marked with pathologies ”Gesangsstimme” and ”Sängerstimme”, which probably contain healthy subjects. However, as it is not clear, we exclude this data unless they are diagnosed with another pathology (see Section 3.1).
@@
From the dataset, we remove 13 corrupted recordings, 40 underage recordings, 17 singers and 335 recordings of patients with multiple recording sessions (except first healthy and first pathological recordings, see Figure 5). See Section 3.1 for rationale and GitHub repository for list of excluded files (misc/ list_of_excluded_files.csv).
@@
When the extraction of f0 was impossible, the feature values for $\overline{f}_{0}$, $\sigma_{\mathbf{f}_{0}}$, $\Delta f_{0}$, jitta, and shim were set to 0 and the binary NaN feature was set to 1 (see Section 3.2).
@@
See Figure 5. More details are in sections 3.1, 3.2, and 3.3.
@@
We utilize multiple ML algorithms for classification (see Table 3) and the k-means SMOTE algorithm for dataset augmentation (see Section 3.3).
@@
All ML algorithms are suitable for multi-dimensional data, that we are dealing with (see Section 3.4).
@@
Information about stratified 10-fold cross-validation and repeated stratified 10-fold cross-validation for the best models is described in the section Section 3.5.
@@
We performed repeated 10-fold cross-validation to estimate the average value of MCC and its corresponding standard deviation. We select the best model according to this MCC. See Section 3.5.
@@
By applying the oversampling algorithm only to the training folds, we aimed to prevent data leakage and ensure that the model’s performance evaluation on the test fold remains unbiased. This approach allowed us to mitigate the adverse effects of class imbalance, resulting in a more robust and reliable predictive model. See Sections 3.3 and 3.5.
@@
For patients with multiple recordings of the same type (either all healthy or all pathological), we retained only the oldest recorded sample. For patients with both healthy and pathological recordings, we selected the oldest sample of each type, resulting in a maximum of two recordings per patient — one healthy and one pathological. We believe this approach minimizes the likelihood of the model learning patient identities, as the patient’s classification remains independent of their identity. See Section 3.1.
@@
All features, except two, are obtained from voice recordings and are widely used in the models for voice pathology detection (see Section 2). We consider the ”AGE” feature legitimate, as other acoustic features depend on the AGE. I.e. there are changes in speaking fundamental frequency with aging [61].
The feature, that we introduced as ”NaN” reflects the fact, that it was not possible to estimate the fundamental frequency for the patient. As the fundamental frequency is considered as one of the dominant features in voice pathology diagnosis, we consider introducing this ”NaN” feature a legitimate approach. Note, that in total, there are 2 females and 6 males in our dataset, that have NaN value of fundamental frequency, all of them suffer from the voice disorder (see Section 3.2).
@@
The choice of metrics, with respect to the class imbalance in the data, is written in the Section 3.5. The claim regarding the best model is based on the Matthews correlation coefficient metric, that is suitable for imbalanced datasets and reflect both successes and errors in the classification.
@@
For each of metrics specified in Section 3.5, we provide also the respective standard deviations that were obtained during the cross-validation procedure which is specified in this section.
@@
First, our model was tested using SVD only. The used database does not fully reflect the general population, in the sense of proportion of healthy / pathological voices. However, at this time, there is no other suitable database that would reflect the general population better. Despite the justification as an only viable source of data, we cannot extrapolate its performance outside of this dataset. Moreover, we limit our research only to individuals that are 18 years old and older. As the data was recorded in a controlled environment, we can assume our models might not be able to perform as well with datasets that are recorded during different conditions. Another noteworthy limitation was the available computational capacity which led to careful decision of the ML algorithms and hyperparameter space we drew from during our work. See Section 4.